OpenStudy (anonymous):

consider the following recurrence relation P(n) = 5P(n-1) - 6P(n-2) if n> 0 and P(1) = 3, P(2)=7 COMPUTE P(N) FOR N=3,4,5,6,7 no idea how to start, Discrete mathematics

OpenStudy (anonymous):

x2-5x+6=0 -> x=-1 , 6 P(n)=a(-1)^n+b6^n P(1)=-a+6b=3 P(2)=a+36b=7 b=5/21 a=30/21-3=-33/21 P(n)=1/21(-33(-1)^n+5.6^n) n=3,4,...